OpenStudy (anonymous):

If A # B = (A x B) / (A + B) a) Calculate 9 # (10 # 10) b) Calculate 8 # (5 # 5)

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

@dan613

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

ok lets to it right \[10\#10=\frac{10\times 10}{10+10}\] is the start so we compute that first

OpenStudy (anonymous):

i get \(\frac{100}{20}=5\)

OpenStudy (anonymous):

it tells you what it means after the equal sign

OpenStudy (anonymous):

it means take the two number, multiply them, then divide by their sum

OpenStudy (anonymous):

so for example \[3\#4=\frac{3\times 4}{3+4}=\frac{12}{7}\]

OpenStudy (anonymous):

ok so continue with the problem

OpenStudy (anonymous):

so lets go back to the top you want \[9\#(10\#10)\]the parentheses mean you have to do that first so step one is compute \(10\#10\)

OpenStudy (anonymous):

yeah i know that

OpenStudy (anonymous):

i get \(10\#10=20\) by the computation aove

OpenStudy (anonymous):

second step is to compute \(9\#20\) as before

OpenStudy (anonymous):

9*20/9+20

OpenStudy (anonymous):

you get \[9\#20=\frac{9\times 20}{9+20}\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so thats a

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@MikeyMaximum

OpenStudy (anonymous):

yes that is A

OpenStudy (anonymous):

youre wrong 9 # (10 # 10) = 9 # (100 / 20) = 9 # 5 = 45 / 14 = 3.21

OpenStudy (anonymous):

i know this because its an online test

OpenStudy (anonymous):

sorry didnt mean to be rude

OpenStudy (anonymous):

oh yes it should have been \(9\#5\) not \(9\#20\) my mistake

OpenStudy (anonymous):

im stills tuck on b i have 1 more try

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

but now that i understand i can do this easily.

OpenStudy (anonymous):

8 # (5 # 5) = 8 # (25 / 10) = 8 # 2.5 = 20 / 10.5 = 1.905 is this right

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

is anyone here

OpenStudy (mikeymaximum):

I'm here now Sorry it took so long :/